paul has twice as many dimes as pennies and 3 times as many nickels as pennies. In all he has $1.80. How many coins of each type does he have?
Answer by Paul(988)

Let dimes be = d
Let nickels be = n
Let pennies = x

n = 3x
d = 2x

Dimes value = 0.1 x 100 =10
Nickel value 0.05 x 100 = 5
Pennie value 0.01 x 100 =1
Total = 1.80 x 100 = 180
10(2x)+5(3x)+(x)=180
20x+15x+x=180
36x=180
x = 5

d = 2(5) = 10
n = 3(5) = 15

Check: 10(10)+5(15)+5 =180
180/100 = $1.80

Hence, there are 5 pennies, 10 dimes and 15 nickels.